OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, 4) and is parallel to y = x - 3.

OpenStudy (tkhunny):

You are given a Point and a hint for the Slope. This suggests very strongly that we should start with the Point-Slope form. State It: (y-y0) = m(x-x0) We have the point (6,4) Substitute: (y-4) = m(x-6) We are almost given the slope. It's a little coded. "parallel to y = x - 3." Well, okay, the slope of that line is "1". Do you see that from y = mx+b? Substitute: (y-4) = (1)(x-6) That's it, really, excepting that we are commanded to put it in Slope-Intercept form. Can you do that?